OpenStudy (anonymous):

You conduct an experiment where you roll a die and record the results. You roll the die 1230 times and the results are shown in the table below: What is the experimental probability of rolling a "4" on your next roll?

OpenStudy (anonymous):

so first you must

OpenStudy (anonymous):

If you can roll a die (6 sided) in such a way that is perfectly random, then there's an equal probability of rolling any number. So, probability of rolling a 4 is 1/6. For experimental probability, you look at what actually happens. How often is a four rolled? If it's rolled, say, 5 times out of 10, then the experimental probability of rolling a 4 is 5/10 = 1/2. experimental probability = # times event occurs/total number of events You can use this to figure out the experimental probability of rolling a 4 next in your problem.